# The Harmonic Exponential Filter for Nonparametric Estimation on Motion Groups

### v1

@@ 5 Beacons 1 Beacon
ATE w/ mode (m) ATE w/ mean (m) NLL (-) ATE w/ mode (m) ATE w/ mean (m) NLL (-)
@@
TABLE III: Quantitative results for the UWB dataset, run II. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). Lower values are better. Note that the EKF is highly uncertain about the ground truth position of the agent during the experiments.  [^3] Complete results at https://montrealrobotics.ca/hef/.
@@ IV-C Doors Dataset
We premapped the environment using gmapping [30] and assumed a bearing-only measurement model that enforces a multimodal distribution. Bearing measurements are computed using the centroid of the doors. We only compute bearing measurements for samples in the free space of our occupancy grid map. Similar to the experiment in Sec. IV-B, we integrate the IMU and wheel encoders to estimate the motion of a Jackal platform. The motion noise was estimated to have a standard deviation of 3.51 cm and 2.50 degrees per step. The ground truth position was estimated using the laser-based SLAM implementation of gmapping. The estimated entropy of the posterior distribution over the ground truth’s pose using gmapping was $2.232$ nats. The layout of the experiment, as well as an image of a door within the dataset is shown in Fig. 6(b).
ATE w/ mode (m) ATE w/ mean (m) NLL (-)
@@
TABLE IV: Quantitative results for the doors dataset. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). We also report the standard deviation of the translation error.
As there is no known correspondence, we employ a greedy data association strategy and compute the measurement likelihood using only the door that is more likely to have generated a given measurement. We present the results of this experiment in Table IV. Our HEF surpasses by a large margin all baselines in this challenging task, showcasing its ability to handle multimodal distributions and the importance of symmetries for robot localization. A snapshot of the filters is shown in Fig. 8.
During these experiments, our proposed approach took an average of 1.14s to run for each timestep. Our approach was implemented in python and ran on an i7-12700H CPU with 16GB of RAM. This was fast enough to be real-time for the door dataset, because the robot would never observe two doors within that time frame.

### v2

@@ 5 Beacons 1 Beacon
ATE w/ mode (m) ATE w/ mean (m) NLP (-) ATE w/ mode (m) ATE w/ mean (m) NLP (-)
@@
TABLE III: Quantitative results for the UWB dataset, run I. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-posterior (NLP). Lower values are better. Note that the EKF is highly uncertain about the ground truth position of the agent during the experiments.  [^3] Complete results at https://montrealrobotics.ca/hef/.
The results of this evaluation are presented in Table III. In all runs and with different beacon configurations, the HEF is comparable to or better than the baseline filters even in the presence of a highly non-Gaussian measurement likelihood. Although the EKF obtained the best ATE among all filters in the first run with one beacon, its NLP is the largest among all filters, indicating that throughout the experiment the filter is highly uncertain about the ground truth position. As all the results of the second run closely resembled those of the first run, they were not included in the table.
@@ IV-C Doors Dataset
We premapped the environment using gmapping [39] and assumed a bearing-only measurement model that enforces a multimodal distribution. Bearing measurements are computed using the centroid of the doors. Bearing measurements are taken only for samples in the free space of the occupancy grid map. As in Sec. IV-B, we integrate the IMU and wheel encoders to estimate the motion of a Jackal platform. The motion noise was estimated to have a standard deviation of 3.51 cm and 2.50 degrees per step. The ground truth position was estimated using the laser-based SLAM implementation of gmapping. The estimated entropy of the posterior distribution over the ground truth’s pose using gmapping was $2.232$ nats. The layout of the experiment, as well as an image of a door within the dataset is shown in Fig. 6(b).
ATE w/ mode (m) ATE w/ mean (m) NLP (-)
@@
TABLE IV: Quantitative results for the doors dataset. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-posterior (NLP). We also report the standard deviation of the translation error.
As there is no known correspondence, we employ a greedy data association strategy, computing the measurement likelihood using only the most likely door for each measurement. We present the results of this experiment in Table IV. Our HEF surpasses by a large margin all baselines in this challenging task, showcasing its ability to handle multimodal distributions and the importance of symmetries for robot localization. A snapshot of the filters is shown in Fig. 7.
During these experiments, our proposed approach took an average of 1.14s to run for each timestep. Our approach was implemented in python and ran on an i7-12700H CPU with 16GB of RAM.